### v1

# Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection

Abstract
Infrared Small Target Detection (IRSTD) system aims to identify small targets in complex backgrounds. Due to the convolution operation in Convolutional Neural Networks (CNNs), applying traditional CNNs to IRSTD presents challenges, since the feature extraction of small targets is often insufficient, resulting in the loss of critical features. To address these issues, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy. First, we propose a selective variable convolution (SVC) module that integrates the benefits of standard convolution, irregular deformable convolution, and multi-rate dilated convolution. This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination of targets from backgrounds. Second, the core component of DCGANet is a two-stage content guided attention module. This module employs two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference. By retaining the most significant responses, this mechanism effectively suppresses false alarms. Additionally, we propose adaptive dynamic feature fusion (ADFF) module to substitute for static feature cascading. This dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, thereby enhancing its ability to discriminate true targets from false alarms. DCGANet has achieved new benchmarks across multiple datasets.
Index Terms:
Infrared small target, selective variable convolution, dynamic content guided attention, adaptive dynamic feature fusion.
I Introduction
Infrared imaging technology, a crucial component of infrared monitoring systems, excels at capturing thermal signatures of targets and demonstrates robust performance in low-light and foggy environments. In military [1], ocean rescue [2], and environmental monitoring [3] applications, targets of interest at long imaging distances often appear as small spots [3]. Consequently, infrared small target detection (IRSTD) is essential for extracting valuable information from complex backgrounds.
Figure 1: The complex background faced by image target detection, the red and green bounding boxes are zoom in area: (a)Dense target of different types such as pedestrians and cars, (b)Target occlusion by other objects such as trees, (c)Strong lighting conditions of sunlight, and (d) Dim lighting conditions during nights.
Despite the critical role of IRSTD in infrared warning systems, as illustrated in Fig. 1, two fundamental challenges remain in the feature extraction of infrared small targets. First, the low resolution of infrared imaging systems results in a paucity of texture and color information. Second, the inherent limitations of imaging small targets occupy an exceedingly small number of pixels. These challenges substantially impede practical performance in complex environments [4]. The primary issues impeding the detection of infrared small targets are as follows:
  1. 1.
    Insufficient features: Infrared small targets often lack distinct structural features due to their low resolution and the absence of texture and color information, making it difficult to differentiate them from complex backgrounds.
  2. 2.
    Low signal-to-noise ratio: The weak infrared radiation signal from small targets, combined with low contrast against background noise, increases the difficulty of detection and recognition.
Addressing these challenges is essential for achieving robust and precise detection in infrared warning systems. It is imperative to suppress background clutter by enhancing the extraction of inherent target features and integrating contextual analysis, which is pivotal to improving IRSTD performance.
Traditional IRSTD methodologies can be categorized into three primary approaches. They are Background Suppression-based (BS-based), Human Visual System-based (HVS-based), and Low Rank Matrix-based (LRM-based) methods. The BS-based methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] utilize specialized filters to extract small targets. The HVS-based methods [16, 17, 18, 19, 20, 21, 22, 23] identify small targets by exploiting local information features. The LRM-based methods [24, 25, 26, 27] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices. However, these traditional methods are limited in their ability to learn from datasets or dynamically adjust parameters in response to changing environmental conditions, leading to suboptimal performance in complex scenarios.
Recent advances in deep learning have yielded substantial improvements for IRSTD. Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [28, 29, 30, 31], contextual information [32, 33, 34, 35, 36], and local contrast [37, 38, 39]. Despite their progress, fundamental challenges in IRSTD remain unaddressed. Specifically, the current limitations are primarily evident in three aspects:
  1. 1.
    Insufficient Feature Extraction: Current convolution methods have shortcomings in capturing details of infrared small targets. Shallow networks have limited receptive fields, which struggle to capture non-local relationship between small targets and background. Meanwhile, deep networks may lose features due to downsampling, leading to inaccurate discrimination between targets and false alarms, thereby affecting the accuracy of detection.
  2. 2.
    Static Attention for Interference Suppression: Traditional convolution and pooling strategies often struggle to effectively separate background and target features, while the saliency of infrared targets may be distributed in specific channels (such as thermal radiation intensity, gradient changes, etc.). However, existing attention mechanisms (such as SE [10], CBAM [34]) model channel relationships statically or at a single scale, causing target information to be masked by background signals, reducing detection sensitivity.
  3. 3.
    Static Feature Fusion: Current feature fusion methods, such as skip connected feature addition or concatenation [40], ignore the differences in various target features at different levels, which can lead to shallow-level noise contamination of deep semantics features, and thus limit the network’s adaptability in dynamic scenes. This further leads to a high false positive rate, affecting the reliability of the detection.
These limitations collectively impede robust detection in real-world scenarios, where targets exhibit dynamic scales and unpredictable noise patterns.
Facing the challenge of IRSTD, existing convolutional neural networks often struggle to effectively address the feature extraction of infrared small targets. To overcome these challenges, there is an imperative need for a novel framework, which is derived from three pivotal observations: (1) Infrared small targets require hierarchical dynamic receptive fields to jointly capture edge sharpness and thermal radiation context; (2) Feature fusion should be selectively guided by high-level semantics to suppress noise propagation; (3) Attention mechanisms must incorporate infrared physical priors and consider input channels according to wavelength sensitivity, while adapting to environmental dynamics. Therefore, our framework should not only effectively extract subtle features of infrared small targets through advanced feature extraction techniques but also dynamically optimize multi-stage feature fusion processes. This framework requires a deep understanding of the background environment, rather than simply relying on local image information near the target.
To address the aforementioned challenges, we propose DGCANet, which consists of three core innovations. First, to improve the recognition of real targets, suppress background noise, and discriminate between them and false alarms, we proposed selective variable convolution (SVC), which consists of multi-branch adaptive kernels. The SVC resolving the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise. Second, a dynamic content guided attention(DCGA) mechanism was proposed. The DGCA first performs local feature extraction and expanded receptive field extraction using parallel ordinary convolution and dilated convolution. To generate multi-scale coarse-grained attention maps, through channel grouping refinement strategy, the input feature channels are grouped according to physical meanings (such as thermal radiation intensity and gradient texture), and captures insensitive features of weak targets through fine attention allocation. Last, we proposed Adaptive Dynamic Feature Fusion (ADFF) to simulate the attention regulation mechanism of the human vision systems(HVS). The integration of ADFF with a wide receptive field significantly improves the distinguish between real targets and false positives.
The main contributions of this paper are listed as follows:
  1. 1.
    We propose DCGANet, a novel framework that unifies feature learning, dynamic content guided attention, and feature fusion in a hierarchical dynamic pipeline. This framework bridges the gap between rigid CNN designs and the dynamic, sparse infrared target, achieving joint optimization of accuracy, robustness, and interpretability.
  2. 2.
    The core of DCGA is to simulate the process of "rapid scanning to fine focusing" in HVS. By using the physical characteristics of infrared imaging, such as the differences in response to targets in different bands, potential target areas were prioritized and located through channel grouping refinement, while irrelevant backgrounds were suppressed.
  3. 3.
    Our SVC combines deformable convolution, multi-dilated convolution, and standard convolution. This architecture enhances the discrimination between real targets and false alarms, and suppresses background noise through its broad receiving domain.
  4. 4.
    The ADFF substantially enhances the main network’s receptive field capacity to differentiate between actual targets and false positives via its sophisticated dynamic fusion technology.
II Related Work
II-A Infrared Small Target Detection
The methodological evolution of infrared small target detection (IRSTD) is divided into model-driven and data-driven paradigms.
II-A1 Model-Driven Infrared Small Target Detection
Model-driven methods formalize IRSTD as anomaly recognition against backgrounds, leveraging three computational paradigms: filters-based, human visual system (HVS) simulations-based, and sparse matrix reconstruction. The filter-based method [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking. For instance, multi-scale Top-Hat [8, 9] filtering algorithm, which combines multi-scale transformation, can improve detection performance. However, when the infrared image contains complex backgrounds, the algorithm will exhibit a lower signal-to-noise ratio (SNR). To further improve the performance, the Human Visual System-Based (HVS-Based) methods [16, 17, 18, 19, 20, 21, 22, 23, 41] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [16], and TLLCM [41]. However, HVS-based methods still struggle to maintain detection stability when faced with high interference from bright edges or points. From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [24, 25, 26, 27, 42], detecting small targets by decomposition low-rank matrices. For instance, the Infrared Patch Image (IPI) [42] model divides infrared images into blocks and recombines them into low-rank sparse matrices. Meanwhile, the reweighted infrared patch tensor (RIPT) [24] was proposed to construct low-rank sparse matrices more stably and faster through the alternating direction method. Overall, although these three categories of methods can achieve good detection results for certain images, their performance dramatically decreases when facing complex backgrounds.
II-A2 Data-Driven Infrared Small Target Detection
Data-driven methods, empowered by deep learning advancements, have emerged as a research frontier in IRSTD. Data-driven methods improve original convolutional neural networks (CNNs) with the prior information of infrared small targets. On the one hand, for the enhancement of local information, the attention local contrast (ALCNet) Net [37] was designed to enhance the local contrast of feature maps. Meanwhile, the dynamic attention transformer (DATNet) [43] was proposed to improve the performance through contextual connectivity. Furthermore, the gated-trans unet (GSTUnet) [29] focuses on the local edge of the small target to improve detection performance. On the other hand, in optimizing network architectures for IRSTD, the dense nested attention (DNANet) [34] employs U-shaped network topology with dense skip connections to achieve feature fusion and target enhancement. Additionally, the background-aware feature exchange net (BAFENet) [44] combines background semantics with feature exchange of targets and backgrounds to achieve high-precision detection.
Although deep learning methods have made substantial progress, they still face challenges such as over-parameterized architectures and inflexible receptive fields [43], impending comprehensive capture of the details of small targets. Excessive reliance on local information [31] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios. DCGANet distinguishes itself through two core innovations:
  1. 1.
    Multi-convolutional spatiotemporal feature extraction: Overcoming CNN rigidity via adaptive kernel configurations across spectral-temporal domains.
  2. 2.
    HVS-mimetic dynamic attention: The DCGA module implements ’coase-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
II-B Vision Attention Mechanisms on IRSTD
Vision attention mechanism in IRSTD enhance target feature expression representation by dynamically prioritizing regions, simultaneously amplifying discriminative target and effectively suppressing noise and false alarms. For instance, the local patch network with global attention net (LPNet) [45] improves the detection accuracy of small targets by combining global and local features. The infrared shape network (ISNet) [31] used Taylor finite difference heuristic blocks and bidirectional attention aggregation modules. In addition, the dynamic gradient attention transformer net (DATransNet) [43] further improves detection performance by introducing a transformer detection module and gradient attention.
Despite these advancements, existing methods still lack robust global feature understanding in IRSTD. In contrast, dynamic content guided attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention. This process refines initial spatial attention maps (SIM) for each channel to focus on the most representative features, enhancing the recognition of suspected small targets. In feature fusion, DCGA adopts a shallow-deep feature fusion strategy to improve information flow between network layers, preserving features and enabling effective gradient backpropagation. This dual optimization not only boosts small target detection accuracy but also strengthens false alarm suppression capabilities.
III Method
III-A Overall Architecture
Figure 2: Overall architecture of the proposed DCGANet, built upon a U-Net-like backbone and integrating the key Selective Variable Convolution (SVC), Dynamic Content Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
The DCGANet architecture, depicted in Fig. 2, is constructed upon a U-Net-like framework and comprises three components: selective variable convolution (SVC), dynamic content guided attention (DCGA), and adaptive dynamic feature fusion (ADFF) modules. Specifically, the encoder path consists of three cascaded layers integrating SVC and DCGA modules in sequence. The SVC replaces standard convolutions to enhance subtle target signatures while expanding receptive fields for augmented context modeling. The DCGA mechanism generates channel-specific spatial importance maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis critical for sub-pixel infrared targets. On the decoder path, two ADFF modules dynamically merge multi-scale encoder features, replacing the original skip-connections and achieving context-aware noise suppression through three-stage fusion.
III-B Selective Variable Convolution
The limited receptive field of traditional convolutional layers in CNNs hinders their ability in IRSTD. To address this, we propose the selective variable convolution (SVC) module, a robust basic block aimed at significantly enhancing the network’s feature learning ability.
As shown in Fig. 3, the SVC combines synergy of the tri-branch features. The first branch employs standard convolution(SConv) to preserve local spatial feature extraction, retaining elementary patterns of small targets and maintaining the stability of fundamental detection performance. The second branch utilizes deformable convolution (DConv), which has been proven effective in edge delineation for irregularly shaped targets, enhancing target features across diverse scales. The third branch, multi-dilation convolution (MDConv), utilizes multi-rate dilated kernels (rates = 2, 4, 8) to extend the receptive field, thereby aggregating multi-scale contexts to suppress structured clutter. The SVC enables accurate extraction of weak target features and significant reduction of background noise.
$X_{s}$ $=\operatorname{SConv}(X_{\text{in}}),$ (1)
$X_{c}$ $=\operatorname{DConv}(X_{\text{in}}),$ (2)
$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}).$ (3)
where $X_{\text{in}}$ is the input feature, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
The output of these branches is fused through channel-wise concatenation and a $1\times 1$ convolution, yielding consolidated representations enriched in both local precision and global discriminability.
$X_{\text{out}}=\operatorname{Conv}_{1\times 1}\left(\operatorname{Concat}\left(X_{s},X_{c},X_{m}\right)\right),$ (4)
where $X_{\text{out}}$ is the output of SVC module, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
By enhancing the extraction of edges and details through a multi-convolution strategy, SVC effectively boosts recognition capability in IRSTD. The expanded receptive field and deformable feature extraction offer substantial advantages in contextual perception and noise suppression, thereby improving the clarity and reliability of feature extraction.
Figure 3: The structure of the SVC: a parallel branching architecture to perform different types of convolution operations on the input feature maps. This design enables the encoding phase to extract features and preserve details at various scales.
III-C Dynamic Content Guided Attention
While the SVC module expands the receptive field and enhances non-local feature extraction, it still struggles to identify the target when dealing with background clutter. Previous attention mechanisms, such as the static channel-wise or spatial weight separation mechanisms of SE [10], CSAM [34], face two key issues. First, there is a lack of information interaction between channel-wise and spatial attention. Second, single-channel spatial weights cannot adapt to the diverse distribution of multiple feature channels and pose a risk of information loss due to the dilution of key target features.
To address these challenges, we propose the dynamic content guided attention(DCGA) module, which uses a coarse-to-fine approach to generate self-attention maps(SIM) for specific channels, effectively simulating the ability of HVS to focus on highlighting features and suppress false alarms. The DCGA is shown in Fig. 4. The first stage includes coarse-grained attention with interaction between channel-wise and spatial attention, while the second stage includes fine-grained calibration with dynamic SIM generation, obtaining a unique SIM for each single-channel input features in a coarse-to-fine manner. We first calculate the corresponding Wc and Ws by following.
III-C1 Stage I: Coarse-Grained Attention with Joint Channel-Space Fusion
The stage include Channel attention, Spatial Attention, and joint attention.
Channel attention: A dual-bottleneck structure extracts global channel responses by aggregating spatial information through global average pooling:
$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right)$ (5)
where $X_{GAP}^{\text{c}}$ denotes channel-wise squeezed features from global average pooling.
Spatial Attention: Spatial saliency is captured by concatenating multi-perspective features from global average pooling (GAP) and global max pooling (GMP):
$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right).$ (6)
where $X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}$ represent spatial descriptors from global average pooling (GAP) and global max pooling (GMP).
Interaction Attention: we fuse $W_{c}$ and $W_{s}$ together using an element-wise addition operation following broadcast rules to obtain a rough SIM $W_{\text{coa}}$.
$W_{\text{coa}}=W_{c}\oplus W_{s},$ (7)
where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$ is rough SIM, $\oplus$ denotes element-wise addition with automatic dimension expansion.
Figure 4: The proposed structure of the DCGA: employs a parallel approach to equally compute the weights of spatial and channel information. It also incorporates random mixing of channel and spatial information to enhance the capability for target segmentation.
III-C2 Stage II: Fine-Grained Calibration with Dynamic SIM Generation
The fine-grained calibration include channel shuffling fusion and dynamic weight generation.
Channel Shuffling Fusion: The raw feature $X\in\mathbb{R}^{C\times H\times W}$ and coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension [46], establishing inter-channel correlations while preserving spatial details.
Dynamic Weight Generation: A gated convolution layer $\mathcal{GC}_{7\times 7}$ followed by sigmoid activation generates channel-specific SIM :
$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right)$ (8)
where $\operatorname{CS}(\cdot)$ denotes the channel shuffling operation. This enables content-adaptive calibration by allowing each channel’s SIM to dynamically adjust based on local feature patterns.
DCGA assigns a unique SIM to each channel, guiding the model to focus on the important areas of each channel. Therefore, more useful information encoded in the features can be emphasized to effectively improve IRSTD.
III-D Adaptive Dynamic Feature Fusion Module
The proposed Adaptive Dynamic Feature Fusion(ADFF) module aims to enhance the network’s ability to distinguish between actual targets and similar false alarms. Specifically, we introduced a dynamic space selection mechanism within the original skip connections, as shown in Fig. 5. This mechanism leverages multi-scale features to enable the network to focus on the areas with the richest spatial contextual information, thereby increasing sensitivity to the target area. The ADFF module is specifically designed to process the feature maps of the encoder-decoder pair in stage i, where the feature maps are represented as:
$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$ (9)
where $C_{i}$ denotes channel number, $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are height and width of feature maps.
Then we applied channel-wise average pooling and max pooling operations to feature maps $X_{i}$and $X^{\prime}_{i}$, represented as $P_{avg}(\cdot)$and $P_{max}(\cdot)$, respectively.
$AD_{i}^{\text{avg}}$ $=\operatorname{AvgPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right),$ (10)
$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right).$ (11)
where $AvgPool$ is average pooling operation, $MaxPool$ is max pooling operation.
These two different pooling operations extract the average and maximum feature descriptors of the space, and the generated spatial pooling features are aggregated into a matrix to facilitate information exchange between different spatial descriptors. Next, we use the convolutional layer $F_{2\to N}(\cdot)$to transform these aggregated features into $N$spatial attention maps. These spatial attention maps can effectively guide the subsequent feature fusion process, highlight salient regions, and suppress redundant background information.
Figure 5: The proposed structure of ADFF: integrates the encoding and decoding stages of the network through a spatially adaptive selection mechanism. It utilizes both maximum and average pooling across channels to achieve precise target extraction.
$AD_{\text{fuse}}=\sigma\left(\operatorname{Conv}_{k\times k}\left(\operatorname{Concat}\left(AD_{i}^{\text{avg}},AD_{i}^{\text{max}}\right)\right)\right),$ (12)
where $\sigma$ denotes the sigmoid function, $\operatorname{Concat}(\cdot)$ is the channel-wise concatenation, and $\operatorname{Conv}_{k\times k}(\cdot)$ represents the convolutional layer (e.g., with a $7\times 7$ kernel, typically outputting a single channel) used to generate the attention map.
The final output $Y_{i}$ of the ADFF module is obtained by first applying the generated attention map $AD_{\text{fuse}}$ to weight the input features $X_{i}$ and $X^{\prime}_{i}$ via element-wise multiplication, and then fusing the results through element-wise addition. This process is formulated as:
$Y_{i}=X_{i}\odot AD_{\text{fuse}}+X^{\prime}_{i}\odot AD_{\text{fuse}}$ (13)
where $\odot$ denotes element-wise multiplication, $Y_{i}$ is the output of ADFF.
The fusion process generates dynamic fusion weights, which can adaptively adjust the importance of each feature map in the final decision. Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby improving the network’s representation ability of IRSTD.
IV Experiment
[…]
table
TABLE I: Configurations for all comparative experiments
Methods(Source & Year) Key parameters configurations
Background Suppression Methods
NWMTH [12]PR’2010 $R_{o}=9,R_{i}=4$
FKRW [47]TGRS’2019 Windows size: 11, $K=4,p=6,\beta=200$
HVS-based Methods
MPCM [17]PR’2016 $N=1,3,\cdots,9$, threshold $k=3$
WLDM [48]TGRS’2016 $L=4,m=2,n=2$, threshold $k=2$
RLCM [49]GRSL’2018 $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
ILCM [50]GRSL’2020 Cell size: $3\times 3$, threshold $k=3$
TLLCM [41]GRSL’2020 Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
GSWLCM [18]GRSL’2022 Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
Low-rank and Sparse Decomposition
IPI [42]TIP’2013 Patch size: 50, sliding step: 10, $\lambda=1/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-7}$
RIPT [24]JSTARS’2017 Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
PSTNN [26]RS’2019 Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
Deep Learning Methods
ACM [32]WACV’2021 Backbone: FPN, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64]
AGPCNet [38]TAES’2023 Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
DNANet [34]TIP’2022 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
ISNet [31]CVPR’2022 Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
UIUNet [33]TIP’2023 Channels: [64, 128, 256, 512], fuse mode: AsymBi
MTUNet [51]TGRS’2023 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
SeRankDet TGRS’2024 Channels:[64, 128, 256, 512, 1024],o=3
SCTransNet [52] TGRS’2024 Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
$\star$ DCGANet (Ours) Channels: [64, 128, 256, 512, 1024]
IV-A Experimental Setup
IV-A1 Datasets
To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST [32], IRSTD-1K [31], and NUDT-SIRST [34]. These datasets demonstrate considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies within real-world operational constraints.
  1. 1.
    NUAA-SIRST: The real long-range infrared single-frame dataset comprises 427 images capturing diverse realistic scenes, such as clouds, cities, and oceans. However, the dataset is relatively small in scale, with only 480 target instances, which may limit its effectiveness for training and evaluating models in scenarios that demand high data diversity.
  2. 2.
    IRSTD-1K: IRSTD-1K is designed to address the challenges of multi-scene cross-domain detection, comprising 1,001 highly diverse images. The complexity of its scene types and target shapes imposes significantly higher demands on the model’s generalization capabilities compared with the other two datasets.
  3. 3.
    NUDT-SIRST: The scene-enhanced dataset, constructed from synthetic data, comprises 1,327 images with complex backgrounds, including cities, fields, bright regions, oceans, and clouds. By employing adaptive target size and intensity blurring functions, it generates diverse targets in the form of point-like, patchy, and extended shapes, yielding distribution characteristics that closely resemble those of real-world datasets.
IV-A2 Evaluation Metrics
To quantitative analyze our DCGANet and other methods, we adopt four standard metrics to evaluate detection performance: Intersection over Union (IoU), normalized IoU (nIoU), Probability of detection (Pd), and False alarm rate (Fa). The detailed description of these metrics are as follows:
  1. 1.
    IoU measures the accuracy of object localization by comparing the overlap between the predicted bounding box and the ground truth. It is defined as the ratio of the intersection area to the union area of the two bounding boxes.
    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
  2. 2.
    nIoU is a standardized adaptation of the IoU metric.
    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
  3. 3.
    Pd quantifies the likelihood of correctly identifying true targets.
    $\text{Pd}=\frac{TP}{TP+FN}$ (16)
  4. 4.
    Fa measures the frequency of incorrect target detections.
    $\text{Fa}=\frac{FP}{FP+TN}$ (17)
where $N$ is the total number of samples, $T$ and $P$ correspond to the number of ground truth and predicted positive pixels, respectively. Additionally, $TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels, respectively.
In addition, the ROC curve illustrates the variation trend of the Probability of detection (Pd) across different False alarm rate (Fa).
IV-A3 Implementation Details
We adopt U-Net as the baseline model, Soft-IoU Loss [53] as the loss function, and AdamW [54] as the optimizer, and incorporating a polynomial learning rate decay strategy to optimize the training process. Additionally, we employ a deep supervision strategy to further enhance model performance. All experiments are conducted using distributed training on a system equipped with four RTX 3090 GPUs. To ensure comprehensive and reliable evaluation, we train independent models on each of the three datasets. The detailed hyperparameter configurations for all comparative methods are summarized in Tab. I. To address the significant differences in dataset distribution and resolution, we customize the hyperparameters for each dataset, with specific configurations provided in Tab. II.
TABLE II: Customized Hyperparameters for Each Dataset
Dataset Epochs Lr Batch Resolution
NUAA-SIRST 500 0.0001 4 512$\times$512
IRSTD-1K 500 0.0001 4 512$\times$512
NUDT-SIRST 500 0.0001 16 256$\times$256
TABLE III: Comparison with Other State-of-the-art methods on four datasets. The $\uparrow$ indicates that the higher the indicator, the better, and the $\downarrow$ indicates that the lower the indicator, the better. The best performance is highlighted in bold, and the second-best is underlined.
Method NUAA-SIRST(Tr=60%) IRSTD-1K(Tr=60%) NUDT-SIRST(Tr=60%)
IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow
                      Background Suppression Methods
NWMTH [12] 15.35 17.00 71.10 55.21 18.31 16.74 49.97 21.06 18.87 16.22 62.84 32.45
FKRW [47] 22.10 28.81 81.41 16.67 10.12 16.83 69.88 24.19 10.04 16.76 72.31 66.90
HVS-based Methods
MPCM [17] 25.64 27.63 68.01 45.07 20.72 23.26 58.78 31.10 27.40 38.95 55.54 11.82
WLDM [48] 5.55 3.37 64.97 21.22 8.16 10.32 54.85 14.79 8.19 11.63 16.30 64.14
RLCM [49] 25.56 27.46 86.34 62.16 16.46 21.83 70.86 67.70 8.21 11.42 70.22 31.67
ILCM [50] 18.42 19.16 38.00 40.63 13.67 11.34 43.32 47.67 5.16 6.48 26.27 88.16
TLLCM [41] 18.20 27.62 82.23 16.18 10.86 18.01 68.95 24.99 13.44 16.98 77.02 69.66
GSWLCM [18] 15.67 13.82 72.48 21.96 12.81 13.46 70.57 13.82 21.08 19.92 69.06 42.78
Low-rank and Sparse Decomposition
IPI [42] 26.67 32.57 83.55 11.97 27.92 30.12 81.27 26.18 25.16 34.64 76.26 43.41
RIPT [24] 11.85 14.91 79.48 22.67 14.71 17.43 79.55 28.35 24.13 33.98 78.54 56.24
PSTNN [26] 22.72 23.50 77.36 29.99 28.83 29.09 82.68 39.48 21.26 24.22 72.78 61.26
Deep Learning Methods
ACM [32] 73.38 74.66 96.06 9.94 60.38 59.83 91.82 21.49 73.86 72.97 97.91 76.22
AGPCNet [38] 77.86 75.23 98.92 34.95 68.30 66.18 94.29 15.55 88.55 88.08 97.40 34.09
DNANet [34] 78.82 79.72 98.63 13.37 69.96 69.35 92.02 23.19 91.32 90.10 97.55 30.71
ISNet[31] 80.83 78.55 98.78 4.93 68.74 61.98 95.80 14.97 88.02 88.76 93.14 28.24
UIUNet [33] 79.76 78.74 98.69 4.41 65.20 64.57 92.33 13.45 88.86 88.11 98.40 23.17
MTUNet [51] 78.41 79.09 98.02 8.66 67.04 68.24 93.18 12.86 87.39 87.34 98.15 29.43
SCTransNet [52] 80.88 79.45 96.48 6.48 71.22 70.11 94.45 13.94 92.39 92.32 97.71 28.76
SeRankDet [55] 81.26 79.66 98.74 2.77 74.45 70.03 95.12 10.65 94.16 93.34 98.71 22.36
$\star$ DCGANet (Ours) 82.03 79.81 98.88 2.99 75.59 72.16 96.64 9.21 94.98 94.46 98.75 20.34
IV-B Comparison With State-of-the-Arts
Our DCGANet method is rigorously compared against state-of-the-arts methods across three datasets using metrics include IoU, nIoU, Pd, and Fa. The BS-based methods NWMTH [12], FKRW [47]. The HVS-based Methods MPCM [17], WLDM [48], RLCM [49], ILCM [50], TLLCM [41], GSWLCM [18]. The LRM-based methods IPI [42], RIPT [24], PSTNN [26]. The deep learning methods ACM [32], AGPCNet [38], DNANet [34], ISNet [31], UIUNet [33], MTUNet [51], SCTransNet [52], SeRankDet [55]. The experimental results summarized in Tab. III, including three critical observations:
1) Our proposed DCGANet exhibits superior performance compared to state-of-the-art methods across multiple datasets. Specifically designed to tackle the challenges posed by low signal-to-noise ratio (SNR) and blurred target boundaries in infrared small target detection, our approach incorporates a multi-modal feature extractor that precisely extracts target features while effectively suppressing background clutter. Additionally, a dedicated feature fusion module is included to aggregate target details from various hierarchical levels. This innovative design substantially enhances detection accuracy and robustness, resulting in significant advancements in the field of infrared small target detection.
2) The experimental results reveal a performance gap exceeding 15% in Intersection over Union (IoU) between Deep Learning-based methods and traditional approaches, highlighting the inherent limitations of model-driven techniques that rely on global sparsity or contrast priors. In contrast, data-driven DL methods demonstrate superior generalization capabilities by leveraging semantic feature extraction and deeper hierarchical representations. This allows them to significantly reduce sensitivity to hyperparameters in unstructured scenarios, resulting in more robust and adaptable performance.
3) On the NUAA-SIRST dataset, DCGANet surpasses SeRankDet across all evaluation metrics except for the False alarm rate (Fa). While SeRankDet shows moderate effectiveness in combining global and local features, its performance significantly deteriorates in cluttered environments. In contrast, DCGANet addresses these limitations by leveraging precise target localization and dynamic feature fusion, resulting in superior pixel-level segmentation accuracy.
Furthermore, Fig. 6 illustrates receiver operating characteristic (ROC) curves for three datasets. The analysis corroborates the numerical results, confirming DCGANet’s stability and superiority over conventional techniques.
(a) NUAA-SIRST
(b) IRSTD-1K
(c) NUDT-SIRST
Figure 6: ROC curves of different methods on three datasets. Our DCGANet achieves the highest Pd at very low Fa.
IV-C Visualization Analysis
To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Fig. 7 - 8). Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarms rates (Fig. 7, yellow boxes) and frequent missed detections (Fig. 8, cyan boxes). In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter effectively, showcasing superior detection robustness.
Gradient-weighted class activation mapping (GradCAM) visualizations for DCGANet (Fig. 9) highlight the focus for each module during the target extraction. The Selective Variable Convolution (SVC) employs diverse convolution forms compared to the baseline UNet, enabling robust feature preservation and edge refinement. The dynamic content guided attention (DCGA) module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets compared to static attention mechanisms. Additionally, the Adaptive Dynamic Feature Fusion (ADFF) module improves edge preservation, achieving sub-pixel accuracy in localization. These results collectively demonstrate that DCGANet’s modular, with each component exhibiting clear and discriminative attention patterns.
Figure 7: Visualization comparison of detection results via different methods on representative images from NUAA-SIRST, IRSTD-1K, and NUDT-SIRST datasets, indicate the land forests, oceans skys, and city building interfere. The red, green, and yellow boxes denote correct detection, false alarms, and missed detections, respectively.
Figure 8: Visualization comparesion the detection results of different methods (The red, green, and yellow boxes denote correct detection, false alarms, and missed detections). 3D mesh plot highlighting the features of small targets in the image.
IV-D Ablation Study
IV-D1 Model-wise performance
Ablation experiments were conducted to evaluate the proposed SVC, DCGA, and ADFF modules, demostrating their individual and integrated impacts on model performance, as shown in Tab. IV. The baseline U-Net was used for comparison analysis.
  1. 1.
    Independent efficacy: Each module yields distinct performance improvements, corroborating its theoretical motivations (e.g. SVC’s 5.50% IoU gain on IRSTD-1K and DCGA’s 7.83% improvement).
  2. 2.
    Convolutional Advancement: Replacing vanilla convolutions with the SVC module elevates IoU by 4.62% on NUAA-SIRST and 5.50% on IRSTD-1K, demonstrating its superiority feature encoding multi-scale target features and suppressing background clutter.
  3. 3.
    Attention Mechanism Superiority:The DCGA outperforms established attention modules (e.g. SE, DAM, and CBAM) by adaptively modeling channel-specific spatial importance maps (SIMs) critical for sub-pixel infrared targets.
  4. 4.
    Contextual Fusion:The ADFF module’s dynamic feature fusion strategy, enhances detection accuracy, achieving a 9.62% improvement over baseline fusion strategies on the IRSTD-1K dataset. This demonstrates the effectiveness of leveraging broader contextual information for target detection in complex infrared scenes.
Furthermore, Fig. 9 presents a qualitative comparison of detection results under different model configurations. The visualizations include original infrared inputs, ground-truth masks, attention maps from the baseline U-Net, and results from the progressively integrated DCGANet, showcasing the incremental benefits of each module in target localization and feature enhancement.
Figure 9: Qualitative comparison of detection results on infrared images. From left to right: original images, predictions by U-Net, U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF/DCGANet, and ground-truth masks. The improvements in target localization and detail are evident with the addition of each module.
TABLE IV: Quantitative Ablation Study on DCGANet (selective variable convolution is marked SVC, dynamic content guided attention is marked DCGA, adaptive dynamic feature fusion module is marked ADFF), The $\uparrow$ indicates that higher values are better.
Strategy Module NUAA-SIRST IRSTD-1K NUDT-SIRST
SVC DCGA ADFF IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
  (a) 75.12 75.35 64.42 62.54 89.81 89.58
(b) 79.74 78.53 71.92 69.51 92.42 92.39
(c) 80.11 78.75 72.25 70.21 93.02 92.88
(d) 81.34 79.01 74.86 72.42 93.98 93.52
(e) $\checkmark$ $\checkmark$ $\checkmark$ 82.03 79.81 75.59 72.16 94.98 94.46
IV-D2 Ablation of SVC module
The SVC module integrates deformable convolutio (DC) and multi-dilation convolution (MDC) alongside Standard Convolution to generate more robust feature representations. Ablation results in Tab. V reveal that while DC alone provides a moderate performance gain, the integration of MDC yields significantly greater improvementds. This demonstrates the superior efficacy of the SVC composite convolutional block, where multi-rate dilated convolutions play a pivotal role in expanding receptive fields and capturing hierarchical contextual information.
These findings underscore the critical importance of broad receptive fields in IRSTD, serving as a key motivator for the design of the ADFF module. By leveraging broader spatial context, the ADFF module is engineered to refine target detection further, enhancing both accuracy and robustness in complex infrared scenes. This synergy between the SVC’s multi-convolutional feature extraction and the ADFF’s contextual fusion highlights the importance of hierarchical receptive field design in overcoming the challenges of low SNR and ambiguous target boundaries.
TABLE V: Ablation experiments of convolution branches of SVC.
Strategy Module NUAA-SIRST IRSTD1K NUDT-SIRST
SC MDC DC IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
(a) 80.35 77.59 72.14 70.12 91.21 91.96
(b) 81.72 78.47 74.01 70.85 93.30 93.88
(c) 80.97 78.22 73.45 70.36 92.64 93.37
(d) 82.03 79.81 75.59 72.16 94.98 94.46
IV-D3 Ablation of DCGA module
Traditional attention modules (SE/DAM/CBAM) fail to address the challenges specific to IRSTD:
  1. 1.
    Under the condition of low SNR, single-channel simulation cannot adequately represent cross-channel variance.
  2. 2.
    Compared with the serial design, the channel shuffling architecture of parallel processing DCGA improves the small target recognition rate.
As shown in Tab. VI, in detection performance, DCGA has significantly improved compared with SE, DAM and CBAM. This shows that DCGA is able to better recalibrate the feature distribution by learning Spatial Importance Maps (SIM) to focus on signal differences between channel-wise, which is particularly critical for enhancing infrared small target features.
TABLE VI: Ablation of DCGA compared with other mainstream attention mechanisms
Strategy NUAA-SIRST IRSTD1K NUDT-SIRST
IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
SE 78.74 77.90 68.13 68.47 90.93 90.98
DAM 79.88 78.01 70.19 69.11 91.29 91.65
CBAM 80.00 77.88 71.20 68.59 92.30 92.04
DCGA 82.03 79.81 75.59 72.16 94.98 94.46
TABLE VII: Ablation study of DCGA. All the experiments are conducted on three datasets.
Method DCGAcas DCGAw/o r DCGAw/o CS DCGA
parallel CA&SA
cascaded CA&SA
refine ✓(remove CS)
NUAA-SIRST IoU 80.12 75.15 78.54 82.03
nIoU 78.71 75.24 77.12 79.81
IRSTD-1K IoU 72.25 64.41 71.34 75.59
nIoU 70.20 62.97 69.04 72.16
NUDT-SIRST IoU 93.04 89.14 91.19 94.48
nIoU 92.78 89.43 91.01 94.46
Further ablation studies on the DCGA design settings are as follows:
  1. 1.
    DCGAcas represents the cascaded connection of channel attention (CA) and spatial attention (SA) in contrast to the parallel mode of DCGA;
  2. 2.
    DCGAw/o r denotes the parallel connection of channel attention (CA) and spatial attention (SA) but without the refinement process;
  3. 3.
    DCGAw/o CS indicates the parallel connection of channel attention (CA) and spatial attention (SA) but excludes the Channel Shuffle (CS) operation in the refinement stage;
  4. 4.
    DCGA is the proposed method that integrates parallel connection of channel attention (CA) and spatial attention (SA), along with a complete refinement process that includes channel shuffle (CS).
The results are presented in Tab. VII, which reveals that model performance slightly declines when the channel shuffle (CS) operation is removed. This result indicates that CS plays a crucial role in facilitating cross-channel information interaction to highlight small target features. Additionally, the channel attention (CA) and spatial attention (SA) mechanisms in DCGA are deployed in parallel, efficiently modeling the salient regions of small targets by generating initial spatial importance maps (SIMs) in a parallel manner. We also experimented with a cascaded design, and the results show that both parallel and cascaded designs exhibit very similar performance in terms of signal-to-noise ratio (SNR) gain. However, when the refinement process is omitted (DCGAw/o CS), performance significantly deteriorates, further confirming the importance of the refinement stage in enhancing the feature representation of infrared small targets.
V Conclusion
In this paper, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adopts a “coarse-to-fine” attention principle to achieve high detection accuracy. First, the selective variable convolution (SVC) module is introduced, integrating standard convolution, deformable convolution, and multi-rate dilated convolution. This module expands the receptive field and enhances non-local features, thereby effectively improving the discrimination between targets and backgrounds. Second, the core component of DCGANet is the dynamic content guided attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference. By retaining the most salient responses, this mechanism suppresses false alarms effectively. Additionally, the adaptive dynamic feature fusion (ADFF) module is proposed to substitute for static feature cascading. The dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, enhancing its capability to discriminate between true targets and false alarms. DCGANet establishes new benchmarks across multiple datasets.